29 percent of the employees of a Fortune 500 company are unhappy. A random sample of 9 employees is selected. What is the probability that the sample contains exactly 3 unhappy employees?
Answer #1

Solution

Given that ,

p = 0.20

1 - p = 1 - 0.20 = 0.80

n = 9

Using binomial probability formula ,

P(X = x) = (n C x) * px * (1 - p)n - x

P(X = 3) = (9 C 3) * (0.20)3 * (0.80)6

= 0.176160

Probability = 0.1762